# Enhancing Joint Motion Prediction for Individuals with Limb Loss Through Model Reprogramming

### 11 Mar 2024 v1

@@ 2 Method
Datasets. The research utilized a public dataset (Hu et al., 2018) for training the gait prediction model on able-bodied individuals, employing time series kinematic sequences from IMUs and goniometers to capture 3D movements and limb positions. The dataset encompassed ten subjects performing various activities, resulting in over 5 million samples for training and 1.3 million for testing. Additionally, data was acquired and processed from three transtibial amputees in a gait laboratory using a Vicon Motion Systems setup with retro-reflective markers for 3D tracking across different locomotion tasks. Gait cycle boundaries were determined using Vicon Nexus software, and joint angles and kinematic data were computed with OpenSim (Delp et al., 2007). Informed consent was obtained from all participants.
@@ 2.3 Refurbish module
We adapt amputee inputs $X_{amp}$ to correct for compensatory motions and asymmetric gait by learning a mapping to a correction template $X_{corr}$, derived from able-bodied data. This is done using a refurbish module $h(.)$ to train on the relationship between $X_{amp}$ and $X_{corr}$ by minimizing two loss functions:
@@
where first term on the R.H.S is the MSE loss between the computed correction $X_{corr}$ (from the template mapping step) and refurbish module predicted correction $\hat{X}_{corr}=h(X_{amp};\Theta_{h})$, the second term on the R.H.S is the MSE loss between the predictions of the pre-trained frozen foundation module $g(.,\Theta^{*}_{g})$ when it is fed with the refurbish module’s outputs and the desired target values of the amputee $y_{amp}$, $\alpha$ and $\beta$ are the factors that influence the effect of the two terms on final loss.

### 12 Mar 2024 v2

@@ 2 Method
Datasets. The research utilized a public dataset (Hu et al., 2018) for training the gait prediction model on able-bodied individuals, employing time series kinematic sequences from Inertial Measurement Units (IMUs) and goniometers to capture 3D movements and limb positions. The dataset encompassed ten subjects performing various activities, resulting in over 5 million samples for training and 1.3 million for testing. Additionally, data was acquired and processed from three transtibial amputees in a gait laboratory using a Vicon Motion Systems setup with retro-reflective markers for 3D tracking across different locomotion tasks. Gait cycle boundaries were determined using Vicon Nexus software, and joint angles and kinematic data were computed with OpenSim (Delp et al., 2007). Informed consent was obtained from all participants.
@@ 2.3 Refurbish module
We adapt amputee inputs $X_{amp}$ to correct for the compensatory motions and asymmetric gait by learning a mapping to a correction template $X_{corr}$, derived from able-bodied data. This is done using a refurbish module $h(.)$ to train on the relationship between $X_{amp}$ and $X_{corr}$ by minimizing two loss functions:
@@
where the first term on the R.H.S is the (mean squared error) MSE loss between the computed correction $X_{corr}$ (from the template mapping step) and refurbish module predicted correction $\hat{X}_{corr}=h(X_{amp};\Theta_{h})$, the second term on the R.H.S is the MSE loss between the predictions of the pre-trained frozen foundation module $g(.,\Theta^{*}_{g})$ when it is fed with the refurbish module’s outputs and the desired target values of the amputee $y_{amp}$, $\alpha$ and $\beta$ are the factors that influence the effect of the two terms on final loss.